OpenStudy (anonymous):

LORD ZARKON! Question for u pls....... 2^x+3^x where x>1000000, find value of x if 2^x+3^x result is divisible by 7

OpenStudy (zarkon):

1,000,005

OpenStudy (zarkon):

x is any number of the form \[3+(n-1)\cdot 6\] where \[n\in\mathbb{N}\]

OpenStudy (anonymous):

OK thts what i got, did u use modulus division

OpenStudy (anonymous):

to see when integer solutions showed up

OpenStudy (zarkon):

yes...I used properties of modular arithmetic.

OpenStudy (anonymous):

is there a neater way then trial an error

OpenStudy (anonymous):

kinda like i was hoping 2^3n=8^n=(7+1)^n

OpenStudy (zarkon):

I din't use trial and error...i derived my formula and then found the smallest n such then 3+(n-1)6 was greater then 1000000 and was divisible by 7

OpenStudy (anonymous):

uh sir, how do u derive the formula without testing so man number

OpenStudy (anonymous):

many*

OpenStudy (anonymous):

cause i saw a pattern

OpenStudy (zarkon):

use 2^3 mod 7 =1 and 3^3 mod 7=6 so x= 3 is a solution 3^6 mod 7 =1 thus 3^9 mod 7=6 from this I get my formula

OpenStudy (anonymous):

uh yh kk

OpenStudy (anonymous):

thnkx once again